I am confused as to solving this linear equation. Should I add them or solve them seperately and how?
2x+7y=11
5x-3y=7.
Thank you.

Answer by ankor@dixie-net.com(22740)

I am confused as to solving this linear equation. Should I add them or solve them seperately and how?
:
This is called a system of equations, two unknowns x & y, two equations that can be added or subtracted to eliminate one of the unknowns.
:
The idea is to mult one or both equations by values that will eliminate one of the unknowns when they are added or subtracted;
:
Looking at this, if we mult the 1st equation by 3 and the 2nd equation by see what we have:
:
2x+7y=11 * 3 =
5x-3y=7. * 7 =
:
6x + 21y = 33
35x -21y = 49
---------------- adding will eliminate y:
41x + 0y = 82
41x = 82
x = 82/41
x = 2
:
Now we know what x is, substiute 2 for x in the 1st equation, find y:
2x + 7y = 11
2(2) + 7y = 11
4 + 7y = 11
7y = 11 - 4
y = 7/7
y = 1
:
You should check the solutions in one of the original equations. we used the 1st one to find y, so the use the 2nd one: 5x - 3y = 7
5(2) - 3(1) = 7
10 - 3 = 7, proves our solution
:
Did this unconfuse you? Any questions